well i`ve been trying this problem for hours "A field has a perimeter of 240 yards. The length is 20 yards less than the width. Find the length and width"?
I cant find the answer.

Answer by checkley71(8403)

perimeter=2*length+2*width
240=2(w-20)+2*w
240=2w-40+2w
4w=240+40
4w=280
w=280/4
w=70 yards then
70-20=50 is the length
proof
240=2*50+2*70
240=100+140
240=240